RHIC Run-6 Dry Run Plan

(based on similar plans for previous RHIC Runs)



RHIC Dry Run (January 23-28)

See also Dry Run - Controls page

Schedule:

Monday 23:                                                       13:30-17:00pm          meeting 3:10pm in MCR

Tuesday 24:        9:00am-12:00pm                   13:30-17:00pm          meeting 9:00am in MCR

Wednesday 25:  9:00am-12:00am   and          13:30-17:00pm           meeting 9:00am   in MCR

Thursday 26:      9:00am-12:00am   and          13:30-17:00pm           meeting 9:00am in  MCR

Friday 27         


(contact Vadim for changes)

 

Mon Jan 23:  Day 1

RHIC operation GPM’s (operations, Vadim)  ramp, magnets, beam current, fill pattern,  etc.                                completed

Logger (RogerL) logger tests                                                                                                                                        completed                                                                 

LogView (Ted) new logview features                                                                                                                            completed

WCM (RogerL) work test                                                                                                                                  completed  

AtR  Line (NickT, Woody) setpoints for pp, ATR manager                                                                                         completed

Permit link (Rob, operations)                                                                                                                                       completed

Abort kickers (Leif, Jian-Lin Mi)                                                                                                                                continues                       

BLAM (Jonathan, Greg) test                                                                                                                                       continues

                        Polarimeter (Haixin, Seth, Sandro) work test                                                                                                             continues

                        Polarimeter emittance measurements (Haixin, Sandro, Anatoly) work test                                                             continues

 

Tue Jan 24:  Day 2 

          Sequencer (Todd, Ted): work test, new features                                                completed  

Elog (Todd, Ila, Greg, Vadim): work test, new features                                                                                               completed

           Plotfill (Todd) test                                                                                                                                                          completed

FDAView (Todd): test of new capabilities                                                                                                                    completed

StartUp (operations, Wolfram): Linux and Sun applications                                                                                       completed          

Configuration (Greg, Vadim) check availability                                                                                                          completed

          AtR Emittance ( NickT, SteveT)  automated measurements, logbook records                                                          completed                    

IPM (RogerC, SteveT) test application features                                                                                                        completed

RhicVacuum Display, Pressure display (DickH, Haixin, Seth): test                                                                         completed

                        e- detectors (SY ): test                                                                                                                                                  completed 

AC Dipole  (Mei) work test                                                                                                                                         completed

 

      Wed Jan 25:  Day 3                        

BPMs (Todd, Rob) Managers, Timing                                                                                                                        continues

                        OrbitDisplay (Vadim) test                                                                                                                                           completed

                        Orbit correction (Vadim, Nikolay) new correction algorithms, ramp correction test                                                continues

Tune meter (Rob, Angelikawork test, high tunes                                                                                                    completed

Transverse injection damper (Rob, Angelika) test                                                                                                    completed

PLL and Feedback (PeteC, Al-DP, Al, Carl):  system tests; tune/coupling feedback                                                continues

SkewMod (Yun) injector, ramping mode                                                                                                                    continues

RhicChromaticity (Steve): work application test for chromaticity and dispersion measurements                            completed 

DQmin (Joanne): work test with Artus, Schottky, PLL                                                                                              completed  

RF applications (TomH.) tests                                                                                                                                    completed

                        BERT (Seth, Greg) test                                                                                                                                              completed

                                    Schottky (Kevin, Al-DP): work test                                                                                                                             completed

                                    Tune display (Dan): test                                                                                                                                               completed

 

 

            Thu Jan 26:    Day 4  

                        Ramping systems (Al, Steve, Nikolay, Todd): pp ramp, WFG Manager,   online model                                        completed

                        RampEditor (Ted) test for new application                                                                                                               completed

                                    Online model (Nikolay,Todd) online-offline model interface, OrbitCalc                                                                  continues

Collimators (Angelika, Wenge): work test, start from sequencer                                                                                Mon

LISA (Angelika, Ted): work test                                                                                                                                  Mon

Gap cleaning (Angelika)  test                                                                                                                                       Mon

RhicInjection (Wolfram, Arlene, Harald): kicker test, application                                                                          completed

IR Bump (Fulvia, Sev): work test, new features                                                                                                       completed

Loss monitors (Mei) RhicLossMonitor, ATR/Injection data acquisition , new BLM’s                                              completed  

 

 

 

 

     Fri Jan 27: Day 5 – back-up session